# Elastic Data Transfer Optimization with Hybrid Reinforcement Learning

### v2

@@ V-C2 Handling Large Files with Parallelism
We designed Dataset B to include files that would take substantial time to transfer if not split into chunks. The per-concurrency bandwidth was throttled to 3 Gbps on the NEWY–CERN path and 1.5 Gbps on the MICH–MASS path. Specifically, the dataset consists of four 55 GB files along with 804 files of size 1 GB. If the large files were to start simultaneously at the beginning, transferring a single 55 GB file with one thread would take approximately 55 GB * 8 / 3 Gbps $\approx$ 147 seconds for the NEWY-CERN path and 55 GB * 8 / 1.5 Gbps $\approx$ 294 seconds for the MICH-MASS path, even under maximum utilization of that thread. To highlight the necessity of parallelism, we deliberately structured the transfer queue such that the large files were not initiated at the same time. Methods with pipelining can efficiently handle these cases by chunking the large files and distributing them across multiple threads. This phenomenon is illustrated for the NEWY-CERN path in Figure 8. The figure shows a steep drop in throughput for Falcon and DRL, both of which lack parallelism. Their throughput curves also exhibit a staircase-like pattern at the tail end. In DRL, one large file was completed before the drop, resulting in steps at 9, 6, and 3 Gbps. In Falcon, the steps occurred at 12 and 6 Gbps, indicating that two large files were completed in close succession. It can be seen that DRL achieves higher throughput than LDMuntil the medium-sized files finish transferring. This occurs because the processing overhead of LDM, as explained in the previous section, outweighs its benefits when sufficient medium files, which are much larger than the maximum possible congestion window, are still available.
This bottleneck directly impacts overall throughput, as shown in Table III and Table IV. In this scenario, LDMclearly outperforms all baselines, achieving approximately 73% higher throughput than DRL on the NEWY–CERN path and 273% higher throughput on the MICH–MASS path. The improvement appears larger on MICH–MASS because the bottleneck files operate at roughly half the per-stream speed (as we bottlenecked) compared to NEWY–CERN, which amplifies the benefits of LDM’s optimizations.
@@ V-C3 Effect of Pipelining on Small Files
Dataset C is designed to demonstrate the effect of pipelining. This dataset consists of 238,770 files with sizes ranging from 100 KB to 1 MB. In the earlier cases, the lack of pipelining did not significantly degrade performance because the number of socket closures and reconnections was relatively small. However, with approximately fifty times more files than Dataset A, the overhead of repeatedly closing and reopening sockets becomes the dominant bottleneck. This effect is evident in the performance comparison shown in Table III and Table IV. Interestingly, Globus outperforms Falcon and DRL in this scenario. Its fixed concurrency setting provides an advantage, since dynamically adjusting concurrency introduces additional socket operations. In contrast, methods that integrate pipelining are able to overcome this bottleneck and clearly outperform the others. Nevertheless, the performance of LDM-GD and LDMis somewhat slower than in other datasets, primarily due to the repeated encoding and decoding of file headers during transfers.
@@ V-D Generalizability Across TCP Congestion Control Algorithms
Most recent work on high-performance data transfers adopts BBR because of its model-driven bandwidth probing and resilience to RTT variations. However, CUBIC remains the default congestion control algorithm in Linux systems and is widely deployed on production servers. To ensure that LDM’s improvements are not specific to BBR, we repeat all experiments on the MICH–MASS testbed using CUBIC and report the results in Table V. The trends remain consistent with the BBR experiments in Table IV.
LDM maintains its relative superiority, outperforming the DRL baseline by approximately 1.1x, 2.3x, 5.9x, and 2.7x across Datasets A, B, C, and D, respectively. For Dataset A, DRL slightly outperformed LDM under BBR because the processing overhead of LDM overshadowed the control-plane bottleneck. Under CUBIC, however, the slower congestion-window growth amplifies control-plane delays, making DRL’s advantage disappear; LDM now leads with 23.76 Gbps versus DRL’s 21.66 Gbps. On Dataset B, where large-file transfers dominate, DRL collapses to 9.16 Gbps due to CUBIC’s conservative window expansion, while LDM reaches 21.09 Gbps through chunk-level parallelism. On Dataset C, pipelining becomes the key factor. Falcon and DRL plateau near 1 Gbps, whereas LDM achieves 7.43 Gbps by continuously issuing file commands and avoiding repeated slow-start resets. For the mixed Dataset D, LDM again delivers the highest throughput at 19.70 Gbps, showing robust adaptation in heterogeneous workloads.
Overall, while CUBIC lowers the achievable throughput ceiling, LDM consistently utilizes a larger fraction of that available bandwidth than all baselines. This confirms that its design generalizes well across both model-based (BBR) and loss-based (CUBIC) congestion control algorithms.
@@
Figure 10: Coexistence of LDM with a fixed-concurrency cotenant.
@@ V-E Fairness Under Multi-tenant Network Conditions
In real scientific environments, multiple data transfers often share the same wide-area link, making fairness an essential requirement for practical deployment. To evaluate how LDM behaves under contention, we run three concurrent transfers on the Fabric NEWY–CERN path and measure their instantaneous and aggregate throughput over time. The results in Figure 9 show that LDM maintains stable sharing behavior and achieves high overall throughput, unlike methods that rely solely on high concurrency. The aggregated throughput for this experiment using Dataset A also matches the values reported in Table III, confirming the consistency of our evaluation. Jain’s Fairness Index (JFI) is a standard metric for quantifying how evenly resources are shared across $n$ flows, defined as:
@@
where $x_{i}$ is the throughput of flow $i$. The index ranges from 0 to 1, with 1 indicating perfectly equal sharing. Figure 9 (c) shows Jain’s Fairness Index over time for one, two, and three concurrent transfers. Fairness remains consistently high (typically above 0.95) across all cases, indicating that LDMallocates bandwidth evenly even as new transfers arrive. Small dips occur at the moments when additional transfers join or depart, but the system quickly returns to a stable high-fairness regime.
To evaluate coexistence with heterogeneous cotenants, we additionally run LDM alongside a static transfer configured with fixed concurrency of 8. As shown in Figure 10, the fixed-concurrency flow maintains a stable throughput of roughly 8 Gbps throughout the experiment, while LDM gradually ramps up and utilizes only the remaining capacity. Even as LDM converges to higher concurrency values, the non-LDM transfer experiences no noticeable slowdown. This indicates that LDM plays fair with heterogeneous co-tenants by harvesting spare bandwidth without exhibiting aggressive behavior. Collectively, these results show that LDM achieves high throughput while preserving fairness and coexistence properties in multi-tenant network environments.
@@ V-F Ablation Study of Transfer Optimizations
We evaluated five configurations: Concurrency only (CC), Pipelining with Concurrency (P-CC), Parallelism with Concurrency (PP-CC), the combined static approach (P-PP-CC), and our fully adaptive solution (LDM), which dynamically optimizes concurrency. The red dashed line in Figures 11 and 12 represents the performance of LDM.
@@
1) Impact on 1 MB Files: For the dataset containing 5120 files of size 1 MB, performance is heavily constrained by control-plane latency and RTT. Since 1 MB is smaller than the maximum per-stream congestion window on both testbeds, solutions that incorporate pipelining are theoretically expected to dominate. On the Fabric testbed (Figure 11a), the CC method reaches only 447 Mbps at concurrency 12, as connection setup overhead dominates the transfer time. Enabling parallelism alone (PP-CC) performs similarly poorly ($\approx$319 Mbps). In contrast, configurations with pipelining (P-CC and P-PP-CC) aggregate small files into long-lived flows, achieving roughly 9.7 Gbps, a 21x speedup over concurrency alone. LDM correctly identifies this requirement during training and converges to an average concurrency of 8.2, achieving 8.2 Gbps. The agent considers additional threads wasteful according to its learned utility function. On Fabric, the gap between the per-stream congestion window and the file size is large, so non-pipelined methods cannot fill the pipe and their bars appear barely visible. On the CloudLab testbed (Figure 12a), this gap is much smaller. As a result, even non-pipelined configurations show visible throughput, although pipelined variants consistently achieve the highest performance. Here as well, LDM reaches the optimal throughput without any manual tuning.
@@
2) Impact on 32 MB Files: This dataset consists of 160 files of size 32 MB, which is larger than the maximum per-stream congestion window on both testbeds. On Fabric, each file is roughly 2.7× larger than the 12 MB per-stream window; on CloudLab, each file is more than 25× larger than the 1.25 MB per-stream window. This means a single file transfer lasts long enough that slow start becomes relatively negligible for CloudLab, and pipelining is not strictly needed to hide control-plane overhead. Likewise, because the dataset contains many independent medium-sized files, parallelism is also unnecessary for both CloudLab and Fabric. For Fabric, pipelining may provide benefits because the per-file transfer time is relatively short compared to the high link capacity. As shown in Figure 11b, the CC configuration on Fabric saturates at approximately 6.4 Gbps. Parallelism alone (PP-CC) performs slightly worse (around 5.4 Gbps) because splitting medium-sized files adds overhead without increasing effective throughput. In contrast, pipelining (P-CC) treats the 160 files as a continuous stream, ensuring that the pipe remains full and reaching roughly 24.2 Gbps. Unlike the 1 MB case, the bars for non-pipelined variants remain visible because each 32 MB transfer is long enough to amortize some overhead. On CloudLab (Figure 12b), all configurations perform nearly identically. Here, the per-file transfer time greatly exceeds command-channel delays, so neither pipelining nor parallelism provides measurable benefit. LDM achieves approximately 24.7 Gbps on Fabric and 975 Mbps on CloudLab, correctly adapting to each environment.
@@
LDM matches these theoretical maxima on both testbeds. On Fabric it converges to roughly 15 threads, which is appropriate given the larger chunk pool and high link speed. On CloudLab it settles near 7 threads, but performance dips slightly due to the unavoidable imbalance caused by the final chunk. Even so, LDM still reaches a competitive throughput for this dataset, demonstrating effective adaptation to file-count limitations and chunk-level parallelism behavior.
Summary: The study confirms that no single static parameter ensures performance across all scenarios. Small and medium files rely heavily on pipelining (p) to overcome latency and BDP constraints, while large sparse datasets require parallelism (pp) to utilize available concurrency. LDM consistently hugs the upper bound of the best-performing static configuration in every scenario, validating its hybrid adaptive design.
@@ V-H Sensitivity to State-Space Design
Figure 13 shows the training rewards for both approaches. The discrete agent learns slowly and exhibits large oscillations, while the continuous agent converges faster and reaches a higher stable reward. These results indicate that directly predicting concurrency provides a smoother optimization landscape and greater sample efficiency. For this reason, LDMadopts the continuous formulation for all experiments.
@@ V-I Discussion
Across all experiments, LDM delivers consistent improvements over existing optimizers. The ablation study shows that pipelining is essential for small files, parallelism is required for large files, and adaptive concurrency aligns the aggregate window with the path BDP. By combining all three, LDM automatically adapts to each dataset without manual tuning.
Generalizability results on CUBIC confirm that LDM’s gains persist even under loss-based congestion control, while the fairness study shows stable sharing in multi-tenant settings. The state-space analysis further demonstrates that continuous actions enable faster and more stable convergence.
By addressing the limitations of static tuning and the training overhead of online learning, LDM offers a comprehensive solution for the "Big Data" transfer challenge.

### v3

@@ V-C2 Handling Large Files with Parallelism
We designed Dataset B to include files that would take substantial time to transfer if not split into chunks. The per-concurrency bandwidth was throttled to 3 Gbps on the NEWY–CERN path and 1.5 Gbps on the MICH–MASS path. Specifically, the dataset consists of four 55 GB files along with 804 files of size 1 GB. If the large files were to start simultaneously at the beginning, transferring a single 55 GB file with one thread would take approximately 55 GB * 8 / 3 Gbps $\approx$ 147 seconds for the NEWY-CERN path and 55 GB * 8 / 1.5 Gbps $\approx$ 294 seconds for the MICH-MASS path, even under maximum utilization of that thread. To highlight the necessity of parallelism, we deliberately structured the transfer queue such that the large files were not initiated at the same time. Methods with pipelining can efficiently handle these cases by chunking the large files and distributing them across multiple threads. This phenomenon is illustrated for the NEWY-CERN path in Figure 8. The figure shows a steep drop in throughput for Falcon and DRL, both of which lack parallelism. Their throughput curves also exhibit a staircase-like pattern at the tail end. In DRL, one large file was completed before the drop, resulting in steps at 9, 6, and 3 Gbps. In Falcon, the steps occurred at 12 and 6 Gbps, indicating that two large files were completed in close succession. It can be seen that DRL achieves higher throughput than EDTuntil the medium-sized files finish transferring. This occurs because the processing overhead of EDT, as explained in the previous section, outweighs its benefits when sufficient medium files, which are much larger than the maximum possible congestion window, are still available.
This bottleneck directly impacts overall throughput, as shown in Table III and Table IV. In this scenario, EDTclearly outperforms all baselines, achieving approximately 73% higher throughput than DRL on the NEWY–CERN path and 273% higher throughput on the MICH–MASS path. The improvement appears larger on MICH–MASS because the bottleneck files operate at roughly half the per-stream speed (as we bottlenecked) compared to NEWY–CERN, which amplifies the benefits of EDT’s optimizations.
@@ V-C3 Effect of Pipelining on Small Files
Dataset C is designed to demonstrate the effect of pipelining. This dataset consists of 238,770 files with sizes ranging from 100 KB to 1 MB. In the earlier cases, the lack of pipelining did not significantly degrade performance because the number of socket closures and reconnections was relatively small. However, with approximately fifty times more files than Dataset A, the overhead of repeatedly closing and reopening sockets becomes the dominant bottleneck. This effect is evident in the performance comparison shown in Table III and Table IV. Interestingly, Globus outperforms Falcon and DRL in this scenario. Its fixed concurrency setting provides an advantage, since dynamically adjusting concurrency introduces additional socket operations. In contrast, methods that integrate pipelining are able to overcome this bottleneck and clearly outperform the others. Nevertheless, the performance of EDT-GD and EDTis somewhat slower than in other datasets, primarily due to the repeated encoding and decoding of file headers during transfers.
@@ V-D Generalizability Across TCP Congestion Control Algorithms
Most recent work on high-performance data transfers adopts BBR because of its model-driven bandwidth probing and resilience to RTT variations. However, CUBIC remains the default congestion control algorithm in Linux systems and is widely deployed on production servers. To ensure that EDT’s improvements are not specific to BBR, we repeat all experiments on the MICH–MASS testbed using CUBIC and report the results in Table V. The trends remain consistent with the BBR experiments in Table IV.
EDT maintains its relative superiority, outperforming the DRL baseline by approximately 1.1x, 2.3x, 5.9x, and 2.7x across Datasets A, B, C, and D, respectively. For Dataset A, DRL slightly outperformed EDT under BBR because the processing overhead of EDT overshadowed the control-plane bottleneck. Under CUBIC, however, the slower congestion-window growth amplifies control-plane delays, making DRL’s advantage disappear; EDT now leads with 23.76 Gbps versus DRL’s 21.66 Gbps. On Dataset B, where large-file transfers dominate, DRL collapses to 9.16 Gbps due to CUBIC’s conservative window expansion, while EDT reaches 21.09 Gbps through chunk-level parallelism. On Dataset C, pipelining becomes the key factor. Falcon and DRL plateau near 1 Gbps, whereas EDT achieves 7.43 Gbps by continuously issuing file commands and avoiding repeated slow-start resets. For the mixed Dataset D, EDT again delivers the highest throughput at 19.70 Gbps, showing robust adaptation in heterogeneous workloads.
Overall, while CUBIC lowers the achievable throughput ceiling, EDT consistently utilizes a larger fraction of that available bandwidth than all baselines. This confirms that its design generalizes well across both model-based (BBR) and loss-based (CUBIC) congestion control algorithms.
@@
Figure 10: Coexistence of EDT with a fixed-concurrency cotenant.
@@ V-E Fairness Under Multi-tenant Network Conditions
In real scientific environments, multiple data transfers often share the same wide-area link, making fairness an essential requirement for practical deployment. To evaluate how EDT behaves under contention, we run three concurrent transfers on the Fabric NEWY–CERN path and measure their instantaneous and aggregate throughput over time. The results in Figure 9 show that EDT maintains stable sharing behavior and achieves high overall throughput, unlike methods that rely solely on high concurrency. The aggregated throughput for this experiment using Dataset A also matches the values reported in Table III, confirming the consistency of our evaluation. Jain’s Fairness Index (JFI) is a standard metric for quantifying how evenly resources are shared across $n$ flows, defined as:
@@
where $x_{i}$ is the throughput of flow $i$. The index ranges from 0 to 1, with 1 indicating perfectly equal sharing. Figure 9 (c) shows Jain’s Fairness Index over time for one, two, and three concurrent transfers. Fairness remains consistently high (typically above 0.95) across all cases, indicating that EDTallocates bandwidth evenly even as new transfers arrive. Small dips occur at the moments when additional transfers join or depart, but the system quickly returns to a stable high-fairness regime.
To evaluate coexistence with heterogeneous cotenants, we additionally run LDM alongside a static transfer configured with fixed concurrency of 8. As shown in Figure 10, the fixed-concurrency flow maintains a stable throughput of roughly 8 Gbps throughout the experiment, while EDT gradually ramps up and utilizes only the remaining capacity. Even as EDT converges to higher concurrency values, the non-EDT transfer experiences no noticeable slowdown. This indicates that EDT plays fair with heterogeneous co-tenants by harvesting spare bandwidth without exhibiting aggressive behavior. Collectively, these results show that LDM achieves high throughput while preserving fairness and coexistence properties in multi-tenant network environments.
@@ V-F Ablation Study of Transfer Optimizations
We evaluated five configurations: Concurrency only (CC), Pipelining with Concurrency (P-CC), Parallelism with Concurrency (PP-CC), the combined static approach (P-PP-CC), and our fully adaptive solution (EDT), which dynamically optimizes concurrency. The red dashed line in Figures 11 and 12 represents the performance of EDT.
@@
1) Impact on 1 MB Files: For the dataset containing 5120 files of size 1 MB, performance is heavily constrained by control-plane latency and RTT. Since 1 MB is smaller than the maximum per-stream congestion window on both testbeds, solutions that incorporate pipelining are theoretically expected to dominate. On the Fabric testbed (Figure 11a), the CC method reaches only 447 Mbps at concurrency 12, as connection setup overhead dominates the transfer time. Enabling parallelism alone (PP-CC) performs similarly poorly ($\approx$319 Mbps). In contrast, configurations with pipelining (P-CC and P-PP-CC) aggregate small files into long-lived flows, achieving roughly 9.7 Gbps, a 21x speedup over concurrency alone. EDT correctly identifies this requirement during training and converges to an average concurrency of 8.2, achieving 8.2 Gbps. The agent considers additional threads wasteful according to its learned utility function. On Fabric, the gap between the per-stream congestion window and the file size is large, so non-pipelined methods cannot fill the pipe and their bars appear barely visible. On the CloudLab testbed (Figure 12a), this gap is much smaller. As a result, even non-pipelined configurations show visible throughput, although pipelined variants consistently achieve the highest performance. Here as well, EDT reaches the optimal throughput without any manual tuning.
@@
2) Impact on 32 MB Files: This dataset consists of 160 files of size 32 MB, which is larger than the maximum per-stream congestion window on both testbeds. On Fabric, each file is roughly 2.7× larger than the 12 MB per-stream window; on CloudLab, each file is more than 25× larger than the 1.25 MB per-stream window. This means a single file transfer lasts long enough that slow start becomes relatively negligible for CloudLab, and pipelining is not strictly needed to hide control-plane overhead. Likewise, because the dataset contains many independent medium-sized files, parallelism is also unnecessary for both CloudLab and Fabric. For Fabric, pipelining may provide benefits because the per-file transfer time is relatively short compared to the high link capacity. As shown in Figure 11b, the CC configuration on Fabric saturates at approximately 6.4 Gbps. Parallelism alone (PP-CC) performs slightly worse (around 5.4 Gbps) because splitting medium-sized files adds overhead without increasing effective throughput. In contrast, pipelining (P-CC) treats the 160 files as a continuous stream, ensuring that the pipe remains full and reaching roughly 24.2 Gbps. Unlike the 1 MB case, the bars for non-pipelined variants remain visible because each 32 MB transfer is long enough to amortize some overhead. On CloudLab (Figure 12b), all configurations perform nearly identically. Here, the per-file transfer time greatly exceeds command-channel delays, so neither pipelining nor parallelism provides measurable benefit. EDT achieves approximately 24.7 Gbps on Fabric and 975 Mbps on CloudLab, correctly adapting to each environment.
@@
EDT matches these theoretical maxima on both testbeds. On Fabric it converges to roughly 15 threads, which is appropriate given the larger chunk pool and high link speed. On CloudLab it settles near 7 threads, but performance dips slightly due to the unavoidable imbalance caused by the final chunk. Even so, EDT still reaches a competitive throughput for this dataset, demonstrating effective adaptation to file-count limitations and chunk-level parallelism behavior.
Summary: The study confirms that no single static parameter ensures performance across all scenarios. Small and medium files rely heavily on pipelining (p) to overcome latency and BDP constraints, while large sparse datasets require parallelism (pp) to utilize available concurrency. EDT consistently hugs the upper bound of the best-performing static configuration in every scenario, validating its hybrid adaptive design.
@@ V-H Sensitivity to State-Space Design
Figure 13 shows the training rewards for both approaches. The discrete agent learns slowly and exhibits large oscillations, while the continuous agent converges faster and reaches a higher stable reward. These results indicate that directly predicting concurrency provides a smoother optimization landscape and greater sample efficiency. For this reason, EDTadopts the continuous formulation for all experiments.
@@ V-I Discussion
Across all experiments, EDT delivers consistent improvements over existing optimizers. The ablation study shows that pipelining is essential for small files, parallelism is required for large files, and adaptive concurrency aligns the aggregate window with the path BDP. By combining all three, EDT automatically adapts to each dataset without manual tuning.
Generalizability results on CUBIC confirm that EDT’s gains persist even under loss-based congestion control, while the fairness study shows stable sharing in multi-tenant settings. The state-space analysis further demonstrates that continuous actions enable faster and more stable convergence.
By addressing the limitations of static tuning and the training overhead of online learning, EDT offers a comprehensive solution for the "Big Data" transfer challenge.